# Chapter 0 Evolution of Triple Stars

### v1

@@ Abstract
  • The complex interplay between stellar evolutionary processes, dissipative processes from stellar interactions, and the gravitational dynamics of triple systems determine their evolution and their byproducts.
@@ 1 Introduction
Triple stellar systems, comprising three gravitationally bound stars, constitute a significant and ubiquitous component of the stellar population. Observational surveys reveal that a substantial fraction of stellar systems, particularly those containing massive stars, originate in triple or higher-order configurations (see Table 1). These systems, encompassing a broad spectrum of stellar masses from low-mass M dwarfs to massive O-type stars, play a pivotal role in various astrophysical phenomena. Their influence extends to the formation of compact and interacting binaries, and they contribute significantly to the genesis of black holes, neutron stars, and progenitors of stellar explosions and mergers (e.g Toonen et al., 2020). Consequently, a thorough understanding of the dynamics and evolution of triple systems and their interplay with dissipative and stellar evolutionary processes is indispensable for a comprehensive understanding of stellar evolution and its implications for astrophysical phenomena.
The dynamics of triple systems exhibit considerably greater complexity compared to binary systems, primarily due to the hierarchical nature of their orbital configurations. Typically, a triple system consists of an inner binary orbited by a more distant tertiary companion (where the tertiary, together with the inner binary center-of-mass are termed the outer-binary in the triple. This hierarchical structure gives rise to a diverse range of dynamical behaviors, spanning from long-term secular evolution [^1] The evolution is considered to be secular if it occurs on timescales much longer than the orbital period of the outer binary of a triple. governed by mechanisms such as von Zeipel-Lidov-Kozai (vZLK) oscillations (see Naoz, 2016, for a review), to quasi-secular [^2] The evolution is considered to be qausi-secular if it occurs on timescales much longer than the orbital period of the inner binary, but could be comparable to the orbit of the outer period. and chaotic interactions that can culminate in stellar collisions or ejections. These multifaceted processes can drive the triple components into intense dynamical and dissipative interactions, encompassing mass loss, mass transfer, and various dissipative processes, thereby profoundly influencing the system’s long-term evolution through close encounters, ejections, and even mergers.
@@ 2 Definition and Structure of Triple Stellar Systems
A triple-stellar system comprises three stars bound by mutual gravitational attraction and engaged in long-term interactions. These systems typically exhibit a hierarchical configuration, wherein two stars form a close binary pair (the inner binary, whose properties are typically denoted by in or the index 1 throughout), while the third star orbits this binary at a significantly greater distance (the outer companion; outer binary, typically denoted by out or the index 2 throughout). This hierarchical arrangement ensures the system’s long-term stability, as the gravitational influence of the outer star is sufficiently weak to avoid disrupting the inner binary on short timescales.
@@
We also introduce additional orbital parameters, used below; these include the argument of the periapsis periastron $\omega$; $\omega_{1}$ and $\omega_{2}$ for the inner and outer orbits ) which is the angle between the ascending node and the periapsis (the point of closest approach between the two bodies in the orbit), measured in the orbital plane. The longitude of the ascending node ($\Omega$; $\Omega_{1}$ and $\Omega_{2}$ for the inner and outer orbits ) is the angle between a reference direction and the ascending node (where the orbit crosses the reference plane), measured in the reference plane. We should also mention the total angular momentum, $j$, and the angular momentum in the z-direction, $j_{z}$; the former is a conserved quantity, and the latter is conserved only in some parameter regimes.
@@ 3 Observed Properties of Triple Systems
Specifically, a diverse range of detection and characterization techniques is necessary to encompass the full spectrum of stars and orbital ranges. These techniques span from spectroscopic analysis and photometry (through potential eclipses) for short-period orbits to interferometric surveys for wider orbits and astrometric measurements for the widest separations (Moe & Di Stefano, 2017; Sana et al., 2014).
@@
Over the past decade and beyond, Tokovinin and Borkovits have identified triple and higher-order multiples among low-mass stars and studied some of their other properties besides their frequency (Tokovinin, 2007, 2014, 2017, 2022). Zinnecker and later Sana et al. have focused on massive stars, demonstrating, as previously mentioned and evident in Table 1, that triples and quadruples are the most prevalent multiplicities among O-stars (Sana et al., 2012, 2014). More recently, in a yet unpublished work, Frost et al. have leveraged combined data from various observational techniques to reveal that the majority of B-stars also exist in triple- or higher-order systems.
@@ 1 Key Orbital and Physical Properties
Frost et al. (in prep.) 6-15 $44$ $>40$
Moe & Di Stefano (2017) 8 – 17 $>57\pm 15$ $61\pm 10$
Sana et al. (2012, 2014) 17 – 50 $>68\pm 18$ $70\pm 11$
@@ 4 Dynamics of Triple Systems
The dynamics of triple stellar systems are rich and complex, encompassing a diverse range of dynamical regimes that govern their long-term evolution. Secular and quasi-secular interactions, in particular, play a critical role in shaping the orbital architecture of hierarchical triples, leading to phenomena such as eccentricity oscillations, inclination oscillations, prograde-retrograde flips, and even chaotic behavior and mergers.
@@
1 Hierarchical Systems: $a_{\mathrm{out}}/a_{\mathrm{in}}\gtrsim 7.5$ $i\lesssim 40^{\circ}$ or $i\gtrsim 140^{\circ}$
@@
2 Secularly Evolving Systems: $a_{\mathrm{out}}/a_{\mathrm{in}}\gtrsim 7.5$ high inclinations; $40^{\circ}\lesssim i\lesssim 140^{\circ}$
@@
3 Quasi-Secular Regime; $3\lesssim a_{\mathrm{out}}/a_{\mathrm{in}}\lesssim 7.5$ high mutual inclinations
@@ 1 Eccentricity excitation by secular and quasi-secular processes
In general, eccentricity excitation by secular processes leads to progressively closer pericenter approaches, to the point where close dissipative interactions affect the evolution, most typically by extracting energy and angular momentum, thereby shrinking and circularizing the orbit. Such processes typically introduce a precession of the orbit, which may quench the secular evolution at some point. Eccentric vZLK, and to a much greater extent, quasi-secular evolution, can drive much faster eccentricity excitations, such that the inner binary may be excited to high eccentricity on a dynamical timescale. Consequently, the binary may experience a sudden jump into a strong interaction phase, potentially even a direct collision, without being previously affected by gradual dissipative processes.
@@
GW emission and inspirals: For compact object inner binaries in triples, for which tides might be far less effective, a sufficiently close approach can introduce large dissipation through gravitational-wave (GW) emission. Similar to the combined secular evolution and tidal friction, combined secular evolution and GW emission can, therefore, drive increased rates of compact binary mergers in triples as suggested in Thompson (2011). In Antonini & Perets (2012), it was shown that rapid secular, and even more so quasi-secular, evolution can lead to sufficiently high eccentricities and not only induce such mergers but drive them to merge while still eccentric, not allowing for the GW dissipation to circularize them before the final merger (see also Antognini et al., 2014). The combined eccentric-vZLK/quasi-secular evolution of compact objects in triples, therefore, drives fast and potentially eccentric GW mergers.

### v2

@@ Abstract
  • The complex interplay between stellar evolutionary processes, dissipative processes from stellar interactions and the gravitational dynamics of triple systems determine their evolution and their byproducts.
@@ 1 Introduction
Triple stellar systems, comprising three gravitationally bound stars, constitute a significant and ubiquitous component of the stellar population. Observational surveys reveal that a substantial fraction of stellar systems, particularly those containing massive stars, originate in triple or higher-order configurations (see Table 1). These systems, encompassing a broad spectrum of stellar masses from low-mass M dwarfs to massive O-type stars, play a pivotal role in various astrophysical phenomena. Their influence extends to the formation of compact and interacting binaries, and they contribute significantly to the genesis of black holes, neutron stars, and progenitors of stellar explosions and mergers (e.g Toonen et al., 2020). Consequently, a thorough understanding of the dynamics and evolution of triple systems, and their interplay with dissipative and stellar evolutionary processes, is indispensable for a comprehensive understanding of stellar evolution and its implications for astrophysical phenomena.
The dynamics of triple systems exhibit considerably greater complexity compared to binary systems, primarily due to the hierarchical nature of their orbital configurations. Typically, a triple system consists of an inner binary orbited by a more distant tertiary companion (where the the teriary, together with the inner binary center-of-mass are termed the outer-binary in the triple. This hierarchical structure gives rise to a diverse range of dynamical behaviors, spanning from long-term secular evolution [^1] The evolution is considered to be secular if it occurs on timescales much longer than the orbital period of the outer binary of a triple. governed by mechanisms such as von Zeipel-Lidov-Kozai (vZLK) oscillations (see Naoz, 2016, for a review), to quasi-secular [^2] The evolution is considered to be qausi-secular if it occurs on timescales much longer than the orbital period of the inner binary, but could be comparable to the orbit of the outer period. and chaotic interactions that can culminate in stellar collisions or ejections. These multifaceted processes can drive the triple components into intense dynamical and dissipative interactions, encompassing mass loss, mass transfer, and various dissipative processes, thereby profoundly influencing the system’s long-term evolution through close encounters, ejections, and even mergers.
@@ 2 Definition and Structure of Triple Stellar Systems
A triple-stellar system comprises three stars bound by mutual gravitational attraction, and engaged in long-term interactions. These systems typically exhibit a hierarchical configuration, wherein two stars form a close binary pair (the inner binary, whose properties are typically denoted by in or the index 1 throughout), while the third star orbits this binary at a significantly greater distance (the outer companion; outer binary, typically denoted by out or the index 2 throughout). This hierarchical arrangement ensures the system’s long-term stability, as the gravitational influence of the outer star is sufficiently weak to avoid disrupting the inner binary on short timescales.
@@
We also introduce additional orbital parameters, used below; these include the argument of the periapsis periastron $\omega$; $\omega_{1}$ and $\omega_{2}$ for the inner and outer orbits ) which is the angle between the ascending node and the periapsis (the point of closest approach between the two bodies in the orbit), measured in the orbital plane. The longitude of the ascending node ($\Omega$; $\Omega_{1}$ and $\Omega_{2}$ for the inner and outer orbits ) is the angle between a reference direction and the ascending node (where the orbit crosses the reference plane), measured in the reference plane. We should also mention the total angular momentum, $j$ and the angular momentum in the z-direction, $j_{z}$; the former is a conserved quantity, and the latter is conserved only in some parameter regimes.
@@ 3 Observed Properties of Triple Systems
Specifically, a diverse range of detection and characterization techniques is necessary to encompass the full spectrum of stars and orbital ranges. These techniques span from spectroscopic analysis and photometry (through potential eclipses) for short-period orbits, to interferometric surveys for wider orbits, and astrometric measurements for the widest separations (Moe & Di Stefano, 2017; Sana et al., 2014).
@@
Over the past decade and beyond, Tokovinin and Borkovits have identified triple and higher-order multiples among low-mass stars and studied some of their other properties, besides their frequency (Tokovinin, 2007, 2014, 2017, 2022). Zinnecker and later Sana et al. have focused on massive stars, demonstrating, as previously mentioned and evident in Table 1, that triples and quadruples are the most prevalent multiplicities among O-stars (Sana et al., 2012, 2014). More recently, in a yet unpublished work, Frost et al. Frost et al. (2025) have leveraged combined data from various observational techniques to reveal that the majority of B-stars also exist in triple- or higher-order systems.
@@ 1 Key Orbital and Physical Properties
Frost et al. (2025) 6-15 $44$ >40>40
Moe & Di Stefano (2017) 8 – 17 $57\pm 15$ $61\pm 10$
Sana et al. (2012, 2014) 17 – 50 $68\pm 18$ $70\pm 11$
@@ 4 Dynamics of Triple Systems
The dynamics of triple stellar systems are rich and complex, encompassing a diverse range of dynamical regimes that govern their long-term evolution. Secular and quasi-secular interactions, in particular, play a critical role in shaping the orbital architecture of hierarchical triples, leading to phenomena such as eccentricity oscillations, inclination oscillations and prograde-retrograde flips, and even chaotic behavior and mergers.
@@
1 Hierarchical Systems: $a_{\mathrm{out}}/a_{\mathrm{in}}\gtrsim 7.5$; $i\lesssim 40^{\circ}$ or $i\gtrsim 140^{\circ}$
@@
2 Secularly Evolving Systems: $a_{\mathrm{out}}/a_{\mathrm{in}}\gtrsim 7.5$; high inclinations; $40^{\circ}\lesssim i\lesssim 140^{\circ}$
@@
3 Quasi-Secular Regime; $3\lesssim a_{\mathrm{out}}/a_{\mathrm{in}}\lesssim 7.5$; high mutual inclinations
@@ 1 Eccentricity excitation by secular and quasi-secular processes
In general, eccentricity excitation by secular processes leads to progressively closer pericenter approaches, to the point where close dissipative interactions affect the evolution, most typically by extracting energy and angular momentum, thereby shrinking and circularizing the orbit. Such processes typically introduce precession of the orbit, which may quench the secular evolution at some point. Eccentric vZLK, and to a much greater extent, quasi-secular evolution, can drive much faster eccentricity excitations, such that the inner binary may be excited to high eccentricity on a dynamical timescale. Consequently, the binary may experience a sudden jump into a strong interaction phase, potentially even a direct collision, without being previously affected by gradual dissipative processes.
@@
GW emission and inspirals: For compact object inner binaries in triples, for which tides might be far less effective, a sufficiently close approach can introduce large dissipation through gravitational-wave (GW) emission. Similarly to the combined secular evolution and tidal friction, combined secular evolution and GW emission can therefore drive increased rates of compact binary mergers in triples as suggested in Thompson (2011). In Antonini & Perets (2012), it was shown that rapid secular, and even more so quasi-secular, evolution can lead to sufficiently high eccentricities and not only induce such mergers but drive them to merge while still eccentric, not allowing for the GW dissipation to circularize them before the final merger (see also Antognini et al., 2014). The combined eccentric-vZLK/quasi-secular evolution of compact objects in triples therefore drives fast and potentially eccentric GW mergers.